OpenStudy (philcua):

need help with homework will fan and medal What is the solution to the equation?

OpenStudy (philcua):

heres the problem i need a solution for x

OpenStudy (kamibug):

\[32 - 6x + 7x = -15 + 28\] First, combine like terms on each side. Can you do that? :)

OpenStudy (philcua):

nope lol this my first homework ,like this

OpenStudy (kamibug):

Like terms have the same variable and exponent. So on the left side the like terms are -6x and 7x -6x + 7x = 1x or just x Get it ?

OpenStudy (philcua):

o ok so the numbers that have the same variable are like terms

OpenStudy (kamibug):

Right. :) So now we have \[32 + x = -15+28\] Can you simplify the right side? :)

OpenStudy (philcua):

ok 13?

OpenStudy (kamibug):

Right ! \[32 + x = 13\] Now to finally solve for x we need to isolate it on one side of the equation. How do you think we can do that? :)

OpenStudy (philcua):

ummm idk moving it to the side

OpenStudy (kamibug):

We can leave x where it is and subtract 32 from both sides to cancel it out on the right. :) Like this...